# Effective Backdoor Mitigation Depends on the Pre-training Objective

### v3

@@ Appendix A Pre-Training Details
We train all the models from scratch on 8 Nvidia A100 GPUs for 64 epochs, an initial learning rate of 0.001 with cosine scheduling and 10000 warmup steps with AdamW optimizer (Loshchilov2017AdamW). The model trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ uses a batch size of 256, whereas the model trained with the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ uses a batch size of 128. We show the change in accuracy and ASR with the pre-training epochs for the model pre-trained on $\mathcal{L}^{pre}_{\text{MMCL}}$ in Figure 13 and for the pre-trained on $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ in Figure 14. We use early-stopping for the model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ and choose the model with the highest accuracy. For $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-training, we choose the model that has the closest accuracy to the $\mathcal{L}^{pre}_{\text{MMCL}}$ pre-trained model. Table 2 shows the accuracy and the ASR of the selected models for both the pre-training objectives.
Appendix B Metrics for the Pre-trained Models
Table 2 shows the Top-1 zero-shot Imagenet Validation set accuracy and the ASR of models that was selected for both the pre-training objectives.
Appendix C Hyperparameter Details
@@
C.1 Cleaning of the Model Pre-trained on CC3M dataset using MMCL
@@
C.2 Cleaning of the Model Pre-trained on CC3M dataset using MMCL and SSL
@@
C.3 Cleaning of the Model Pre-trained on CC6M dataset using MMCL
@@
C.4 Cleaning of the Model Pre-trained on CC6M dataset using MMCL and SSL
@@
C.5 Cleaning of the Model Pre-trained on the CC3M dataset under Non-Ideal Conditions
@@
C.6 Cleaning of the Model Pre-trained on the CC6M dataset under Non-Ideal Conditions
@@
C.7 Cleaning of the Model Pre-trained on the CC6M dataset using an Objective distinct from Pre-training
@@
C.8 Cleaning of the Model Pre-trained on the CC6M dataset using Heavy Regularization
@@
C.9 Cleaning of the Model Pre-trained on the CC6M dataset using Shrink and Perturb
@@
C.10 Cleaning of the Model Pre-trained on the CC6M dataset using a Larger Cleaning Dataset
@@
C.11 Cleaning of the Model Pre-trained on the CC6M dataset with More Finetuning Epochs
@@
C.12 Cleaning of the Model Pre-trained on the CC6M dataset using a Larger Weights for SSL Term
@@
Appendix D CC3M Results
D.1 Findings for the Models trained on the CC3M Dataset
Figure 7: Top-1 zero-shot Imagenet validation set accuracy v/s the ASR, measured at the end of each finetuning epoch for the models pre-trained on the CC3M dataset. The finetuning is done with each of the three aforementioned losses. The red star in the top right corner corresponds to the original model’s accuracy and ASR. For a successful cleaning, there should be models that maintain the original model’s accuracy while having a low ASR (indicated by the red circle). Takeaway: CleanCLIP successfully cleans the model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ (left), while it fails for the model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ (right).
Figure 7 shows the scatter plot of the Top-1 zero-shot Imagenet-1K validation set accuracy and the ASR of the models at the end of each finetuning epoch for the models pre-trained on the CC3M dataset. We observe that:
@@ D.1 Findings for the Models trained on the CC3M Dataset
D.2 Findings for the Models trained on the CC3M Dataset when Cleaning under Non-ideal Conditions
@@
Figure 8: Scatter plot of the Top-1 zero-shot Imagenet validation set accuracy v/s the ASR during the finetuning process for the models pre-trained on the CC3M dataset. The finetuning is done with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. We measure accuracy and ASR at the end of each epoch. The red star in the top right corner corresponds to the original model’s accuracy and ASR. For a successful cleaning, there should be models that maintain the original model’s accuracy while having a low ASR (indicated by the red circle). Takeaway: Even having 5 poisons in the cleaning dataset (i.e. 0.005% of the dataset, which is 10$\times$ cleaner than the pre-training data) hurts the cleaning process for both pre-training objectives, and $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained models are hurt worse.
@@
Figure 8 shows the scatter plot of the Top-1 zero-shot Imagenet validation set accuracy and the ASR at the end of each finetuning epoch for the models pre-trained on the CC3M dataset. We only show the models finetuned with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. We observe that having just 5 poisoned datapoints in the finetuning dataset severely lessens the effectiveness of CleanCLIP for both the pre-training objectives. However for the models pre-trained with just $\mathcal{L}^{pre}_{\text{MMCL}}$, we found cleaned models that maintain the original model’s accuracy and get around 30-50% ASR. On the other hand, for the models pre-trained with the stronger objective $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$, having just 5 poisoned examples renders the cleaning procedure completely ineffective. No model has an ASR lower than 90% for this model.
Appendix E Cleaning with a Larger Cleaning Dataset
@@
In this experiment we doubled the size of the finetuning data to 200K, which is guaranteed to be clean, and finetuned the pre-trained model on this dataset using 14 learning rates for 20 epochs. Figure 9 shows the scatter plot of the Top-1 Imagenet validation set zero-shot accuracy and the ASR at the end of each finetuning epoch. Even after doubling the finetuning datasize, CleanCLIP is ineffective for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model, as it loses about 90% of the original accuracy to get an ASR $\leq$ 5%. See Appendix C for the hyperparameters we explored for this experiment.
@@
Figure 9: The scatter plot of the Top-1 zero-shot Imagenet validation set accuracy v/s the ASR at the end of each finetuning epoch for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M dataset. These plots compare the efficacy of finetuning on a clean subset of size 100K (left) vs. 200K (right) datapoints. Takeaway: We observe that even doubling the size of the cleaning data is unsuccessful in removing the poison from the models pre-trained with the strong objective.
Appendix F Cleaning with More Finetuning Epochs

### v4

@@ Appendix A Pre-Training Details
We train all the models on 8 Nvidia A100 GPUs for 64 epochs. We use an initial learning rate of 0.001 for the models trained from scratch, and for the models where poison is induced by finetuning from a pre-trained checkpoint, we use an initial learning rate of $75e-$ 7. We use cosine scheduling and 10000 warmup steps with AdamW optimizer (Loshchilov & Hutter, 2017) for training. The model trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ uses a batch size of 256, whereas the model trained with the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ uses a batch size of 128.
We show the change in accuracy and ASR with the training epochs for the model trained from scratch with BadNet attack using $\mathcal{L}^{pre}_{\text{MMCL}}$ in Figure 17 and using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ in Figure 18. We use early-stopping for the model trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ and choose the model with the highest accuracy. For $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-training, we choose the model that has the closest accuracy to the $\mathcal{L}^{pre}_{\text{MMCL}}$ trained model. Table 3 shows the accuracy and the ASR for all the models we select in this paper for poison removal.
Appendix B Performance of the Pre-trained Models
Table 3 shows the Top-1 zero-shot Imagenet Validation set accuracy and the ASR of models that were selected for both the pre-training objectives.
Appendix C Templates for Text-Embedding Computation
a bad photo of a {class}.’, a photo of many {class}.’, a sculpture of a {class}.’, a photo of the hard to see {class}.’, a low resolution photo of the {class}.’, a rendering of a {class}.’, graffiti of a {class}.’, a bad photo of the {class}.’, a cropped photo of the {class}.’, a tattoo of a {class}.’, the embroidered {class}.’, a photo of a hard to see {class}.’, a bright photo of a {class}.’, a photo of a clean {class}.’, a photo of a dirty {class}.’, a dark photo of the {class}.’, a drawing of a {class}.’, a photo of my {class}.’, the plastic {class}.’, a photo of the cool {class}.’, a close-up photo of a {class}.’, a black and white photo of the {class}.’, a painting of the {class}.’, a painting of a {class}.’, a pixelated photo of the {class}.’, a sculpture of the {class}.’, a bright photo of the {class}.’, a cropped photo of a {class}.’, a plastic {class}.’, a photo of the dirty {class}.’, a jpeg corrupted photo of a {class}.’, a blurry photo of the {class}.’, a photo of the {class}.’, a good photo of the {class}.’, a rendering of the {class}.’, a {class} in a video game.’, a photo of one {class}.’, a doodle of a {class}.’, a close-up photo of the {class}.’, a photo of a {class}.’, the origami {class}.’, the {class} in a video game.’, a sketch of a {class}.’, a doodle of the {class}.’, a origami {class}.’, a low resolution photo of a {class}.’, the toy {class}.’, a rendition of the {class}.’, a photo of the clean {class}.’, a photo of a large {class}.’, a rendition of a {class}.’, a photo of a nice {class}.’, a photo of a weird {class}.’, a blurry photo of a {class}.’, a cartoon {class}.’, art of a {class}.’, a sketch of the {class}.’, a embroidered {class}.’, a pixelated photo of a {class}.’, itap of the {class}.’, a jpeg corrupted photo of the {class}.’, a good photo of a {class}.’, a plushie {class}.’, a photo of the nice {class}.’, a photo of the small {class}.’, a photo of the weird {class}.’, the cartoon {class}.’, art of the {class}.’, a drawing of the {class}.’, a photo of the large {class}.’, a black and white photo of a {class}.’, the plushie {class}.’, a dark photo of a {class}.’, itap of a {class}.’, graffiti of the {class}.’, a toy {class}.’, itap of my {class}.’, a photo of a cool {class}.’, a photo of a small {class}.’, a tattoo of the {class}.’
Appendix D Hyperparameter Details
@@
D.1 Cleaning of the Model Pre-trained on CC3M dataset using MMCL
@@
D.2 Cleaning of the Model Pre-trained on CC3M dataset using MMCL and SSL
@@
D.3 Cleaning of the Model Pre-trained on CC6M dataset using MMCL
@@
D.4 Cleaning of the Model Pre-trained on CC6M dataset using MMCL and SSL
@@
D.5 Cleaning of the Model where Poison is induced via Finetuning with MMCL on CC6M dataset
Cleaning Epochs: 20
Learning rates (69 values): {5e-05, 4.9e-05, 4.8e-05, 4.7e-05, 4.6e-05, 4.5e-05, 4.4e-05, 4.3e-05, 4.25e-05, 4.2e-05, 4.1e-05, 4e-05, 3.9e-05, 3.8e-05, 3.75e-05, 3.7e-05, 3.6e-05, 3.5e-05, 3.4e-05, 3.3e-05, 3.2e-05, 3.1e-05, 3e-05, 2.9e-05, 2.8e-05, 2.7e-05, 2.6e-05, 2.5e-05, 2.4e-05, 2.3e-05, 2.2e-05, 2.1e-05, 2e-05, 1.9e-05, 1.8e-05, 1.7e-05, 1.6e-05, 1.5e-05, 1.4e-05, 1.3e-05, 1.2e-05, 1.1e-05, 1e-05, 9e-06, 8e-06, 7e-06, 6e-06, 5e-06, 4.9e-06, 4.75e-06, 4.5e-06, 4.25e-06, 4e-06, 3.75e-06, 3.5e-06, 3.25e-06, 3e-06, 2.75e-06, 2.5e-06, 2.25e-06, 2e-06, 1.75e-06, 1.5e-06, 1.25e-06, 1e-06, 5e-07, 1e-07, 5e-08, 1e-08}
MMCL weight: 1
SSL weight: 1
Size of the Cleaning Dataset: 1,00,000
D.6 Cleaning of the Model where Poison is induced via Finetuning with MMCL + SSL on CC6M dataset
Cleaning Epochs: 20
Learning rates (85 values): {0.005, 0.001, 0.0005, 0.0001, 5e-05, 5e-05, 4.9e-05, 4.8e-05, 4.75e-05, 4.7e-05, 4.6e-05, 4.5e-05, 4.5e-05, 4.4e-05, 4.3e-05, 4.25e-05, 4.2e-05, 4.1e-05, 4e-05, 4e-05, 3.9e-05, 3.8e-05, 3.75e-05, 3.7e-05, 3.6e-05, 3.5e-05, 3.5e-05, 3.4e-05, 3.3e-05, 3.25e-05, 3.2e-05, 3.1e-05, 3e-05, 3e-05, 2.9e-05, 2.8e-05, 2.75e-05, 2.7e-05, 2.6e-05, 2.5e-05, 2.5e-05, 2.4e-05, 2.3e-05, 2.25e-05, 2.2e-05, 2.1e-05, 2e-05, 2e-05, 1.9e-05, 1.8e-05, 1.75e-05, 1.7e-05, 1.6e-05, 1.5e-05, 1.5e-05, 1.4e-05, 1.4e-05, 1.3e-05, 1.3e-05, 1.2e-05, 1.2e-05, 1.1e-05, 1.1e-05, 1e-05, 1e-05, 9e-06, 9e-06, 8e-06, 8e-06, 7e-06, 7e-06, 6e-06, 6e-06, 5e-06, 5e-06, 1e-06, 1e-06, 5e-07, 5e-07, 1e-07, 1e-07, 5e-08, 5e-08, 1e-08, 1e-08}
MMCL weight: 1
SSL weight: 1
Size of the Cleaning Dataset: 1,00,000
D.7 Cleaning of the Model with Label Consistent Poisoning trained with MMCL on CC6M dataset
Cleaning Epochs: 20
Learning rates (44 values): {5e-05, 4.75e-05, 4.25e-05, 4e-05, 3.8e-05, 3.75e-05, 3.7e-05, 3.6e-05, 3.5e-05, 3.4e-05, 3.3e-05, 3.2e-05, 3.1e-05, 3e-05, 2.9e-05, 2.8e-05, 2.7e-05, 2.6e-05, 2.5e-05, 2.4e-05, 2.3e-05, 2.2e-05, 2.1e-05, 2e-05, 1.9e-05, 1.8e-05, 1.7e-05, 1.6e-05, 1.5e-05, 1.4e-05, 1.3e-05, 1.2e-05, 1.1e-05, 1e-05, 9e-06, 8e-06, 7e-06, 6e-06, 5e-06, 1e-06, 5e-07, 1e-07, 5e-08, 1e-08}
MMCL weight: 1
SSL weight: 1
Size of the Cleaning Dataset: 1,00,000
D.8 Cleaning of the Model with Label Consistent Poisoning trained with MMCL + SSL on CC6M dataset
Cleaning Epochs: 20
Learning rates (71 values): {0.007, 0.006, 0.005, 0.004, 0.003, 0.002, 0.001, 0.0009, 0.0008, 0.0007, 0.0006, 0.0005, 0.0004, 0.0003, 0.0002, 0.00018, 0.00017, 0.00016, 0.00014, 0.00013, 0.00012, 0.00011, 0.0001, 9e-05, 8e-05, 7e-05, 6e-05, 5e-05, 4.75e-05, 4.25e-05, 4e-05, 3.8e-05, 3.75e-05, 3.7e-05, 3.6e-05, 3.5e-05, 3.4e-05, 3.3e-05, 3.2e-05, 3.1e-05, 3e-05, 2.9e-05, 2.8e-05, 2.7e-05, 2.6e-05, 2.5e-05, 2.4e-05, 2.3e-05, 2.2e-05, 2.1e-05, 2e-05, 1.9e-05, 1.8e-05, 1.7e-05, 1.6e-05, 1.5e-05, 1.4e-05, 1.3e-05, 1.2e-05, 1.1e-05, 1e-05, 9e-06, 8e-06, 7e-06, 6e-06, 5e-06, 1e-06, 5e-07, 1e-07, 5e-08, 1e-08}
MMCL weight: 1
SSL weight: 1
Size of the Cleaning Dataset: 1,00,000
D.9 Cleaning of the Model Pre-trained on the CC3M dataset under Non-Ideal Conditions
@@
D.10 Cleaning of the Model Pre-trained on the CC6M dataset under Non-Ideal Conditions
@@
D.11 Cleaning of the Model Pre-trained on the CC6M dataset using an Objective distinct from Pre-training
@@
D.12 Cleaning of the Model Pre-trained on the CC6M dataset using Heavy Regularization
@@
D.13 Cleaning of the Model Pre-trained on the CC6M dataset using Shrink and Perturb
@@
D.14 Cleaning of the Model Pre-trained on the CC6M dataset using a Larger Cleaning Dataset
@@
D.15 Cleaning of the Model Pre-trained on the CC6M dataset with More Finetuning Epochs
@@
D.16 Cleaning of the Model Pre-trained on the CC6M dataset using a Larger Weights for SSL Term
@@
Appendix E CC3M Results
E.1 Findings for the Models trained on the CC3M Dataset
Refer to caption 20% drop
Figure 10: Top-1 zero-shot Imagenet validation set accuracy vs. the ASR, measured at the end of each cleaning epoch for the models pre-trained on the CC3M dataset. The finetuning is done with each of the three losses as mentioned above. The red star in the top right corner (encircled in the black circle) corresponds to the original model’s accuracy and ASR. For a successful cleaning, there should be models that maintain the original model’s accuracy while having a low ASR (indicated by the red circle). Takeaway: CleanCLIP successfully cleans the model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ (left), while it fails for the model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ (right).
Figure 10 shows the scatter plot of the Top-1 zero-shot Imagenet-1K validation set accuracy and the ASR of the models at the end of each cleaning epoch for the models pre-trained on the CC3M dataset. We observe that:
@@ E.1 Findings for the Models trained on the CC3M Dataset
E.2 Findings for the Models trained on the CC3M Dataset when Cleaning under Non-ideal Conditions
@@
Figure 11: Scatter plot of the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR during the cleaning process for the models pre-trained on the CC3M dataset. The finetuning is done with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. We measure accuracy and ASR at the end of each epoch. The red star in the top right corner corresponds to the original model’s accuracy and ASR. For a successful cleaning, there should be models that maintain the original model’s accuracy while having a low ASR (indicated by the red circle). Takeaway: Even having 5 poisons in the cleaning dataset (i.e. 0.005% of the dataset, which is 10$\times$ cleaner than the pre-training data) hurts the cleaning process for both pre-training objectives, and $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained models are hurt worse.
@@
Figure 11 shows the scatter plot of the Top-1 zero-shot Imagenet validation set accuracy and the ASR at the end of each cleaning epoch for the models pre-trained on the CC3M dataset. We only show the models finetuned with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. We observe that having just 5 poisoned datapoints in the finetuning dataset severely lessens the effectiveness of CleanCLIP for both the pre-training objectives. However, for the models pre-trained with just $\mathcal{L}^{pre}_{\text{MMCL}}$, we found cleaned models that maintain the original model’s accuracy and get around 30-50% ASR. On the other hand, for the models pre-trained with the stronger objective $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$, having just 5 poisoned examples renders the cleaning procedure completely ineffective. No model has an ASR lower than 90% for this model.
Appendix F Effectiveness of CleanCLIP when Poison is Induced using a Different Backdoor
Refer to caption 16% drop
Figure 12: Scatter plot of the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR at the end of each cleaning process for the models poisoned with label consistent poison. The finetuning is done with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. We measure the accuracy and ASR at the end of each finetuning epoch. The red star in the top right corner (encircled in the black circle) corresponds to the original model’s accuracy and ASR. For a successful clean, there should be models that maintain the original model’s accuracy while having a low ASR (indicated by the red circle). Takeaway: CleanCLIP is much more effective for the model trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ (left) than for the model trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ (right).
In this experiment, we poison models using a different kind of backdoor: label consistent backdoor. In this backdoor, we add a trigger patch to an image whose caption contains the adversary chosen label, in our experiment “banana”. Therefore, in this case, the adversary does not need to change the labels of the poisoned datapoints. Similar to the previous experiments, we trained two models, one using $\mathcal{L}^{pre}_{\text{MMCL}}$ and the other using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ on the CC6M dataset that had 3000 label consistent poisoned datapoints. We train the models from scratch using a starting learning rate of $31e-$ 3 using cosine scheduling with 10,000 warmup steps with AdamW optimizer.
After training the models, we chose two models that had similar accuracy and cleaned them using CleanCLIP, i.e., finetuned them using a clean dataset of 100K image-text pairs using $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$, using several learning rates (refer to Appendix D for hyperparameter details). We measure the Top-1 Imagenet validation set accuracy and ASR for the models at the end of each cleaning epoch and plot the scatter plot for the two metrics in Figure 12.
We observe that similar to BadNet poisoning, CleanCLIP is much more effective for the model trained with the simpler $\mathcal{L}^{pre}_{\text{MMCL}}$ objective. The model trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ lose 16% accuracy compared to the model trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ that gains 10% accuracy, to obtain a model with a low ASR ($\leq$ 5%).
Appendix G Cleaning with a Larger Cleaning Dataset
@@
In this experiment, we doubled the size of the finetuning data to 200K, which is guaranteed to be clean, and finetuned the pre-trained model on this dataset using 14 learning rates for 20 epochs. Figure 13 shows the scatter plot of the Top-1 Imagenet validation set zero-shot accuracy and the ASR at the end of each cleaning epoch. Even after doubling the finetuning data size, CleanCLIP is ineffective for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model, as it loses about 90% of the original accuracy to get an ASR $\leq$ 5%. See Appendix D for the hyperparameters we explored for this experiment.
@@
Figure 13: The scatter plot of the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR at the end of each cleaning epoch for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M dataset. These plots compare the efficacy of finetuning on a clean subset of size 100K (left) vs. 200K (right) datapoints. Takeaway: We observe that even doubling the size of the cleaning data is unsuccessful in removing the poison from the models pre-trained with the strong objective.
Appendix H Cleaning with More Finetuning Epochs